### v1

@@ 3.3 The Osmosis Echo
Repeated application of $\bm{P}$ yields the state transition matrix $\bm{S}$ of the steady state. Since its eigenvalues are given by powers of the eigenvalues of $\bm{P}$, we can conclude that it has a single nonzero eigenvalue $\lambda_{1}=1$. Thus, $\bm{S}$ is a rank-$1$ matrix, given by the outer product of $\bm{q}_{1}$ and the corresponding row of $\bm{Q}^{-1}$. Since $\bm{P}^{\top}=\left(\bm{Q}\bm{\Lambda}\bm{Q}^{-1}\right)^{\top}=(\bm{Q}^{-\top}\bm{\Sigma}\bm{Q}^{T})$, the column of $\bm{Q}^{-\top}$ corresponding to $\lambda_{1}$ (which is the corresponding eigenvector of $\bm{P}^{\top}$) is constant, so the corresponding row of $\bm{Q}^{-1}$ is also constant. It follows that all columns of the steady-state state transition matrix are the same. They are given by the eigenvector $\bm{q}_{1}$ of $\bm{P}$, which is the steady-state solution of the process.
@@ 3.4 The Optic Flow Echo
Figure 3 shows an example of a flow field computed with the Horn–Schunck model [13]. The flow field displays the estimated motion between two frames of a test sequence from the Middlebury dataset [53]. We use a colour-code [54] for the visualisation of the flow field, where the colours encode different flow directions.
Refer to caption Refer to caption Refer to caption
(a) frame 12 (b) frame 13 (c) computed flow
Figure 3: Estimation of the optic flow field (colour-coded) using the Horn–Schunck method ($\alpha=10000$) on frame 12 and frame 13 of the Urban test sequence from the Middlebury flow data set [53]. See e.g. [54] for the used colour code.
@@ 5.1.2 Nonlinear Diffusion Echoes
(a) original (b) filtered (c) echo $(120,72)$ (d) echo $(168,136)$
@@ 5.1.5 Optic Flow Echoes
(a) ref. frame (b) HS flow (c) HS echo zoom
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
(d) normal flow (e) NE flow (f) NE echo zoom
Figure 11: Computed colour-coded optic flow of a rectangle moving one pixel to the right and example echo. The white pixel in the normal flow and computed flow marks the echo location. The flow is computed with the Horn–Schunck (HS) model ($\alpha=10$) and the Nagel–Enkelmann (NE) method ($\alpha=1$, $\lambda=0.5$). We visualise echoes from the central pixel of the left vertical line of the normal flow. Since the normal flow in this pixel only has a horizontal direction, we only consider the echo for the horizontal component. The echoes have a negligible vertical component, so instead of colour coding them, we logarithmically rescale the horizontal component and visualise them in greyscale, marking the echo location in red. This enhances the visibility of the details. The normal flow information is locally transported into the rectangle. However, the Horn–Schunck model also propagates flow information outside of the rectangle, leading to a blurry result.
@@ 5.2 Echo Compression
To evaluate the results, we proceed twofold. First, we compute the error $\lVert\bm{S}-\widehat{\bm{S}}\rVert_{F}$ between the state transition matrix and the approximation. Since a calculation of this would require us to compute $\bm{S}$ explicitly, we instead apply Hutchinson’s trace estimator [65], using that $\lVert\bm{A}\rVert_{F}^{2}=\operatorname{trace}(\bm{A}^{T}\bm{A})$. In addition to the quantitative evaluation, we visualise a few reconstructed source echoes to get a better intuition about the reconstruction quality.
@@
Table 1: Errors for RSVD-based compression. Estimated [65] Frobenius norm of the error depending on the fraction of used singular values.
percentage of singular values $0.5\,\%$ $1.25\,\%$ $2.5\,\%$ $5.0\,\%$
NLD (PM) $2.198$ $0.666$ $0.074$ $0.012$
NLD (We) $37.550$ $30.949$ $16.279$ $0.008$
EED (PM) $0.015$ $0.012$ $0.012$ $0.012$
@@ 5.2.1 An Extension for Rapidly Decaying Diffusivities
Table 2: Errors for isotropic nonlinear diffusion with the Weickert diffusivity using the proposed exclusion mechanism with $\epsilon\in\{0.0,0.05,0.1\}$. Estimated [65] Frobenius norm of the error depending on the fraction of used singular values.
percentage of singular values $0.5\,\%$ $1.25\,\%$ $2.5\,\%$ $5.0\,\%$
$\epsilon=0.0$ $37.550$ $30.949$ $16.279$ $\bm{0.008}$
$\epsilon=0.05$ $22.378$ $12.168$ $\bm{0.745}$ $0.745$
$\epsilon=0.1$ $\bm{16.617}$ $\bm{4.528}$ $1.596$ $1.596$
@@
It should be further noted that the approach even reduces the storage cost slightly, since each excluded point reduces the size of the singular vectors, while only requiring the storage of two integers.

### v2

@@ 3.3 The Osmosis Echo
Repeated application of $\bm{P}$ yields the state transition matrix $\bm{S}$ of the steady state. Since its eigenvalues are given by powers of the eigenvalues of $\bm{P}$, we can conclude that it has a single nonzero eigenvalue $\lambda_{1}=1$. Thus, $\bm{S}$ is a rank-$1$ matrix, given by the outer product of $\bm{q}_{1}$ and the corresponding row of $\bm{Q}^{-1}$. Since $\bm{P}^{\top}=\left(\bm{Q}\bm{\Lambda}\bm{Q}^{-1}\right)^{\top}=(\bm{Q}^{-\top}\bm{\Lambda}\bm{Q}^{T})$, the column of $\bm{Q}^{-\top}$ corresponding to $\lambda_{1}$ (which is the corresponding eigenvector of $\bm{P}^{\top}$) is constant, so the corresponding row of $\bm{Q}^{-1}$ is also constant. It follows that all columns of the steady-state state transition matrix are the same. They are given by the eigenvector $\bm{q}_{1}$ of $\bm{P}$, which is the steady-state solution of the process.
@@ 3.4 The Optic Flow Echo
Figure 3 shows an example of a flow field computed with the Horn–Schunck model [13]. The flow field displays the estimated motion between two frames of a test sequence from the Middlebury dataset [53].
Refer to caption Refer to caption Refer to caption Refer to caption
(a) frame 12 (b) frame 13 (c) computed flow (d) colour code
Figure 3: Estimation of the optic flow field (colour-coded) using the Horn–Schunck method ($\alpha=10000$) on frame 12 and frame 13 of the Urban test sequence from the Middlebury flow data set [53] and the used colour code.
@@ 5.1.2 Nonlinear Diffusion Echoes
(a) original (b) filtered (c) echo in $(120,72)$ (d) echo in $(168,136)$
@@ 5.1.5 Optic Flow Echoes
(a) ref. frame (b) HS flow (c) zoom into (a) (d) HS echo zoom
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
(e) normal flow (f) NE flow (g) zoom into (e) (h) NE echo zoom
Figure 11: Computed colour-coded optic flow of a square moving one pixel to the right and example echo. The flow is computed with the Horn–Schunck (HS) model ($\alpha=10$) and the Nagel–Enkelmann (NE) method ($\alpha=1$, $\lambda=0.5$). We use the same colour code as in Figure 3, so the ground truth motion would be a red square. We visualise echoes from the central pixel of the left vertical line of the normal flow. This location is marked in white in the normal flow and the computed flow fields and in red in the corresponding echoes. Since the normal flow in this pixel only has a horizontal direction (see (e)), we only consider the echoes for the horizontal component. These echoes have a negligible vertical component, so instead of colour coding them, we logarithmically rescale the horizontal component and visualise them in greyscale. This enhances the visibility of the details. We display zooms of the echoes which correspond to the zooms in (c) and (g). The echoes (d) and (h) visualise how the normal flow information is locally transported to the inside of the square. However, (d) shows that the Horn–Schunck model also propagates flow information to the outside of the square, which leads to the blurry result from (b).
@@ 5.2 Echo Compression
To evaluate the results, we proceed twofold. First, we compute the error $\lVert\bm{S}-\widehat{\bm{S}}\rVert_{F}$ between the state transition matrix and the approximation. Since a calculation of this would require us to compute $\bm{S}$ explicitly, we instead apply Hutchinson’s trace estimator [64] with Gaussian random vectors [65], using that $\lVert\bm{A}\rVert_{F}^{2}=\operatorname{trace}(\bm{A}^{T}\bm{A})$. We select the number of samples based on the results from Avron and Toledo [66]. Since we do not need the highest precision, we select the number so that the median relative approximation error is at most 10%. However, experiments in [66] suggest that we can expect substantially higher accuracy in practice. Nevertheless, we also complement this approximation-based quality criterion with a deterministic one: We compute the MSE between the filtered image $\bm{u}$ and the reconstruction $\widehat{\bm{S}}\bm{f}$ based on the approximated state transition matrix. In addition to the quantitative evaluation, we visualise a few reconstructed source echoes to get a better intuition about the reconstruction quality.
@@
Table 1: Errors for RSVD-based compression. Estimated [64] Frobenius norm of the error and MSE of the reconstruction of the filtered image depending on the fraction of used singular values.
percentage of Frobenius Norm / MSE
singular values $0.5\,\%$ $1.25\,\%$ $2.5\,\%$ $5.0\,\%$
NLD (PM) $2.198$ / $0.84810$ $0.666$ / $0.02864$ $0.074$ / $0.00034$ $0.012$ / $0.00010$
NLD (We) $37.55$ / $6451.53$ $30.95$ / $3637.93$ $16.28$ / $534.286$ $0.008$ / $0.00291$
EED (PM) $0.015$ / $0.00012$ $0.012$ / $0.00010$ $0.012$ / $0.00010$ $0.012$ / $0.00010$
@@ 5.2.1 An Extension for Rapidly Decaying Diffusivities
To find the pixels that should be excluded, we first need to compute the corresponding echoes, increasing the cost of the encoding phase. However, we can reduce this additional cost by decreasing the number of potential exclusion candidates for which the echoes must be computed. To this end, recall that the state transition matrix $\bm{S}$ for diffusion processes has nonnegative entries and unit row sums (see Section 3.1.1), and that the initial image $\bm{f}$ is also nonnegative. Then it follows from 5 that if the diagonal element $s_{ii}>1-\epsilon$, we must have $(1-\epsilon)\,f_{i}+\epsilon\,f_{min}<u_{i}<(1-\epsilon)\,f_{i}+\epsilon\,f_{max}$ for the corresponding pixel $u_{i}$ of the filtered image. If any of these inequalities is violated, we can conclude that $s_{ii}\leq 1-\epsilon$ and the corresponding echo will not be excluded. However, fulfilling the inequalities does not guarantee $s_{ii}>1-\epsilon$, so we need to manually check by explicitly computing the echo.
@@
Table 2: Errors for isotropic nonlinear diffusion with the Weickert diffusivity using the proposed exclusion mechanism with $\epsilon\in\{0.0,0.05,0.1\}$. Estimated [64] Frobenius norm of the error and MSE of the reconstruction of the filtered image depending on the fraction of used singular values.
percentage of Frobenius Norm / MSE
singular values $0.5\,\%$ $1.25\,\%$ $2.5\,\%$ $5.0\,\%$
$\epsilon=0.0$ $37.55$ / $6451.53$ $30.95$ / $3637.93$ $16.28$ / $534.286$ $\bm{0.008}$ / $\bm{0.00291}$
$\epsilon=0.05$ $22.38$ / $2736.83$ $12.17$ / $269.923$ $\bm{0.745}$ / $\bm{0.01346}$ $0.745$ / $0.01358$
$\epsilon=0.1$ $\bm{16.62}$ / $\bm{1282.20}$ $\bm{4.528}$ / $\bm{9.19979}$ $1.596$ / $0.05048$ $1.596$ / $0.05048$
@@
Although the required precomputation of the echoes for exclusion increases the cost of the encoding phase, it should be further noted that the exclusion approach even reduces the storage cost slightly, since each excluded point reduces the size of the singular vectors while only requiring the storage of two integers.